### v1

@@ 1 Introduction
Despite the previous efforts, three significant obstacles remain: (i) Many stochastic control problems in economics are finite-horizon time-inhomogeneous problems, which may be more difficult than the related infinite horizon problem, as the optimal control policy at different time period is different. (ii) Due to the curse of dimensionality, it is generally difficult to numerically solve stochastic control problems in high dimensions and for problems with complicated stochastic dynamics. (iii) If the utility function in the control problem is not time-separable, then such a problem may not have the Bellman equation.
To overcome these difficulties, we propose a machine learning algorithm, the Monotonic Monte Carlo Control (MMCC) algorithm, to solve high-dimensional, finite time horizon, and time-inhomogeneous stochastic control problems without using dynamic programming principles. The MMCC algorithm can be implemented using a deep neural network representation of the policy functions where the parameters of the neural networks can be learned by stochastic gradient descent approach. In each round of training, the algorithm first generates sample paths of the states and controls by Monte Carlo simulation, and then updates the control policies in each time period sequentially in a backward direction. Therefore, The MMCC algorithm has a monotonicity of performance improvement in each iteration step, leading to good convergence properties. The algorithm does not require the Bellman equation and does not require the utility function to be time-separable, and allows general stochastic dynamics of the evolution of states. We demonstrate the effectiveness of the MMCC algorithm by solving various high-dimensional (e.g., over 100-dimensional) stochastic control problems including portfolio selection under a stochastic volatility model, multi-sector stochastic growth, and optimal control under a dynamic stochastic integration of climate and economy model with 8-dimensional state vectors and 600 time periods.
@@ 1.1 Literature Review
For infinite-horizon and heterogeneous-agent (HA) problems, grid-free methods based on simulation are proposed. Maliar, Maliar and Winant (2021) demonstrate the effectiveness of their grid-free framework by solving two representative-agent models and one heterogeneous-agent model. Additionally, Han, Yang and E (2021) employ both supervised and unsupervised learning using two neural networks in a grid-free manner, with unsupervised learning applied to policy function approximation and supervised learning to value function approximation. They essentially aim to maximize lifetime reward and then evaluate their method based on Bellman-equation error. Moreover, Hall-Hoffarth (2023) minimizes the error of equilibrium conditions to solve a heterogeneous-agent New Keynesian (HANK) model.
@@
For finite-horizon problems, existing literature focuses on solving overlapping generations models by grid-free methods. For example, Duarte, Fonseca, Goodman and Parker (2021) solve a life-cycle model by maximizing lifetime reward; Azinovic, Gaegauf and Scheidegger (2022) and Azinovic and Žemlička (2023) solve an overlapping generations model by minimizing the error of equilibrium conditions. In particular, Azinovic and Žemlička (2023) introduce a market clear layer to enforce the economic constraints.
@@
This paper complements the existing literature in four aspects: (i) The proposed MMCC algorithm is a grid point-free (i.e., simulation-based) algorithm, which avoids the curse of dimensionality and enables us to handle high dimensional problems (e.g., 100 dimensions). In this regard, our algorithm is closely related to the problem of American option pricing using simulation (see, e.g., Longstaff and Schwartz (2001), Tsitsiklis and Van Roy (2001), and Broadie and Glasserman (1997, 2004), Glasserman (2004, Ch. 8)). (ii) The MMCC algorithm differs from existing grid point-free methods in the way network parameters are updated. In each round of iteration, the MMCC algorithm updates the neural networks of the control policies at different time periods sequentially in a backward manner. In contrast, all neural network parameters are updated simultaneously in existing grid point-free methods. (iii) The MMCC algorithm has a monotonicity of performance improvement at each iteration, while many existing algorithms do not have such a property. (iv) The MMCC algorithm does not use the Euler equation or Bellman equation; in contrast, many numerical algorithms and grid point-free algorithms in the literature rely on the Euler equation, Bellman equation, or their approximation. (v) The MMCC algorithm can solve problems with time-inseparable utility functions, which may not have Bellman equations.
@@
Besides the economic literature, there is a large literature in applied mathematics and applied probability about stochastic control. In particular, Yong and Zhou (1999) and Flemming and Soner (2006) provide an in-depth discussion on continuous time stochastic control problems and their applications. Kushner and Dupuis (2001) give an excellent survey of numerical methods for solving continuous time stochastic control problems by using Markov chains. There have also been many studies on the numerical solutions to continuous time stochastic control problems in mathematical finance. [^1] See, e.g., Zhang (2004), Bouchard and Touzi (2004), Chassagneux (2014), Chassagneux and Richou (2016), Crisan, Manolarakis and Touzi (2010), Gobet, Lemor and Warin (2005), Gobet and Turkedjiev (2016, 2017), Henry-Labordere, Tan and Touzi (2014), Henry-Labordère, Oudjane, Tan, Touzi and Warin (2019), Kharroubi, Langrené and Pham (2015), Kharroubi, Langrené and Pham (2014), and Guo, Zhang and Zhuo (2015). Most of these studies focus on particular stochastic processes, e.g., discretized diffusion processes or Lévy processes, but our MMCC algorithm can be applied to general stochastic processes. Moreover, our method is a simulation-based method, suitable for high-dimensional problems.
@@
Approximate dynamic programming (ADP) has been developed [^2] ADP has also evolved under the name of reinforcement learning in computer science (see, e.g., Sutton and Barto (1998)). for dealing with three sources of curses of dimensionality: high dimensionality of state space, control policy space, and random shock space; see the books by Powell (2011) and Bertsekas (2012). ADP algorithms can be broadly classified into two categories: value iteration and policy iteration. [^3] Many ADP algorithms focus on infinite time horizon problems where the optimal value function and policy are stationary. In contrast, our MMCC algorithm focuses on finite time horizon problems where neither the optimal value function nor the optimal policy is stationary. Most ADP algorithms are value iteration algorithms, which approximate the value function by employing the Bellman equation. [^4] Value function iteration is closely related to the duality approach for stochastic dynamic programming; see, e.g., Brown, Smith and Sun (2010), Brown and Smith (2014), Brown and Haugh (2017), and Chen, Ma and Yu (2016). As an alternative, a policy iteration algorithm keeps track of the policy instead of the value function. At each period, a value function is calculated based on a policy estimated previously and then improved within the policy space. However, the value iteration and policy iteration ADP algorithms may not have a monotonic improvement of the value function at each iteration. The MMCC algorithm is related to but is fundamentally different from the policy iteration ADP algorithms mainly in that: (i) The MMCC algorithm does not use the Bellman equation; (ii) The MMCC algorithm has a monotonic improvement of the value function at each iteration; (iii) The MMCC algorithm can be applied to general control problems in which the objective functions may not be time-separable.

### v2

@@ 1 Introduction
Despite the previous efforts, three significant obstacles remain: (i) Many stochastic control problems in economics are finite-horizon time-inhomogeneous problems, which may be more difficult than the related infinite horizon problem, as the optimal control policies at different time periods are different. (ii) Due to the curse of dimensionality, it is generally difficult to numerically solve stochastic control problems in high dimensions and for problems with complicated stochastic dynamics. (iii) If the utility function in the control problem is not time-separable, then such a problem may not have the Bellman equation.
To overcome these difficulties, we propose a machine learning algorithm, the Monotonic Monte Carlo Control (MMCC) algorithm, to solve high-dimensional, finite time horizon, and time-inhomogeneous stochastic control problems without using dynamic programming principles. The MMCC algorithm can be implemented using a deep neural network representation of the policy functions where the parameters of the neural networks can be learned by stochastic gradient descent approach. In each round of training, the algorithm first generates sample paths of the states and controls by Monte Carlo simulation, and then updates the control policies in each time period sequentially in a backward direction. Therefore, The MMCC algorithm has a monotonicity of performance improvement in each iteration step, leading to good convergence properties. The algorithm does not require the Bellman equation, does not require the utility function to be time-separable, and allows general stochastic dynamics of the evolution of states. We demonstrate the effectiveness of the MMCC algorithm by solving various high-dimensional (e.g., over 100-dimensional) stochastic control problems including portfolio selection under a stochastic volatility model, multi-sector stochastic growth, and optimal control under a dynamic stochastic integration of climate and economy model with 8-dimensional state vectors and 600 time periods.
@@ 1.1 Literature Review
For infinite-horizon and heterogeneous-agent (HA) problems, grid-free methods based on simulation are proposed. Maliar, Maliar and Winant (2021) demonstrate the effectiveness of their grid-free framework by solving two representative-agent models and one heterogeneous-agent model (Krusell and Smith, 1998). Additionally, Han, Yang and E (2022) employ both supervised and unsupervised learning using two neural networks in a grid-free manner, with unsupervised learning applied to policy function approximation and supervised learning to value function approximation. They essentially aim to maximize lifetime reward and then evaluate their method based on Bellman-equation error. Moreover, Hall-Hoffarth (2023) minimizes the error of equilibrium conditions to solve a heterogeneous-agent New Keynesian (HANK) model. In the case of continuous-time infinite-horizon HA problems, Huang (2023) focuses on the probabilistic formulation of economic models and uses deep learning to solve the system of forward-backward stochastic differential equations corresponding to the model. Subsequently, Huang (2024) employs the probabilistic approach to solve a continuous-time version of Krusell and Smith (1998) and an asset pricing model with search-and-bargaining frictions.
@@
For finite-horizon problems, existing literature focuses on solving overlapping generations models by grid-free methods. For example, Duarte, Fonseca, Goodman and Parker (2021) solve a life-cycle model by maximizing lifetime reward; Azinovic, Gaegauf and Scheidegger (2022) and Azinovic and Jan Žemlička (2023) solve an overlapping generations model by minimizing the error of equilibrium conditions. In particular, Azinovic and Jan Žemlička (2023) introduce a market clear layer to enforce the economic constraints.
@@
This paper complements the existing literature in four aspects: (i) The proposed MMCC algorithm is a grid point-free (i.e., simulation-based) algorithm, which avoids the curse of dimensionality and enables us to handle high dimensional problems (e.g., 100 dimensions). In this regard, our algorithm is closely related to the problem of American option pricing using simulation (see, e.g., Longstaff and Schwartz (2001), Tsitsiklis and Van Roy (2001), and Broadie and Glasserman (1997, 2004), Glasserman (2004, Ch. 8)). (ii) The MMCC algorithm differs from existing grid point-free methods in how network parameters update. In each round of iteration, the MMCC algorithm updates the neural networks of the control policies at different time periods sequentially in a backward manner. In contrast, all neural network parameters are updated simultaneously in existing grid point-free methods. (iii) The MMCC algorithm has a monotonicity of performance improvement at each iteration, while many existing algorithms do not have such a property. (iv) The MMCC algorithm does not use the Euler equation or Bellman equation; in contrast, many numerical algorithms and grid point-free algorithms in the literature rely on the Euler equation, Bellman equation, or their approximation. (v) The MMCC algorithm can solve problems with time-inseparable utility functions, which may not have Bellman equations.
@@
Besides the economic literature, there is a large body of literature on applied mathematics and applied probability in stochastic control. In particular, Yong and Zhou (1999) and Flemming and Soner (2006) provide an in-depth discussion on continuous time stochastic control problems and their applications. Kushner and Dupuis (2001) give an excellent survey of numerical methods for solving continuous time stochastic control problems by using Markov chains. There have also been many studies on the numerical solutions to continuous time stochastic control problems in mathematical finance. [^1] See, e.g., Zhang (2004), Bouchard and Touzi (2004), Chassagneux (2014), Chassagneux and Richou (2016), Crisan, Manolarakis and Touzi (2010), Gobet, Lemor and Warin (2005), Gobet and Turkedjiev (2016, 2017), Henry-Labordere, Tan and Touzi (2014), Henry-Labordère, Oudjane, Tan, Touzi and Warin (2019), Kharroubi, Langrené and Pham (2015), Kharroubi, Langrené and Pham (2014), and Guo, Zhang and Zhuo (2015). Most of these studies focus on particular stochastic processes, e.g., discretized diffusion processes or Lévy processes, but our MMCC algorithm can be applied to general stochastic processes. Moreover, our method is a simulation-based method, suitable for high-dimensional problems.
@@
Approximate dynamic programming (ADP) has been developed [^2] ADP has also evolved under the name of reinforcement learning in computer science (see, e.g., Sutton and Barto (1998)). for dealing with three sources of curses of dimensionality: high dimensionality of state space, control policy space, and random shock space; see the books by Powell (2011) and Bertsekas (2012). ADP algorithms can be broadly classified into two categories: value iteration and policy iteration. [^3] Many ADP algorithms focus on infinite time horizon problems where the optimal value function and policy are stationary. In contrast, our MMCC algorithm focuses on finite time horizon problems where neither the optimal value function nor the optimal policy is stationary. Most ADP algorithms are value iteration algorithms, which approximate the value function by employing the Bellman equation. [^4] Value function iteration is closely related to the duality approach for stochastic dynamic programming; see, e.g., Brown, Smith and Sun (2010), Brown and Smith (2014), Brown and Haugh (2017), and Chen, Ma, Liu and Yu (2024). As an alternative, a policy iteration algorithm keeps track of the policy instead of the value function. At each period, a value function is calculated based on a policy estimated previously and then improved within the policy space. However, the value iteration and policy iteration ADP algorithms may not have a monotonic improvement of the value function at each iteration. The MMCC algorithm is related to but is fundamentally different from the policy iteration ADP algorithms mainly in that: (i) The MMCC algorithm does not use the Bellman equation; (ii) The MMCC algorithm has a monotonic improvement of the value function at each iteration; (iii) The MMCC algorithm can be applied to general control problems in which the objective functions may not be time-separable.